OpenStudy (anonymous):

x''+x' = 100sin5t Find a particular solution xp(t) of the given differential equation using the complex function method

OpenStudy (anonymous):

@Neemo do u find anything ?

OpenStudy (anonymous):

yeaaaah ! I have connection problem...equation editor isn't working well...W8 pleaase ! okk ?

OpenStudy (anonymous):

ok of course

OpenStudy (anonymous):

I can guide you if you want to ! my equation editor is not working anymore !!!! :(

OpenStudy (anonymous):

seriously :/

OpenStudy (anonymous):

i want of course but my english is not perfect. i may have a problem

OpenStudy (anonymous):

we will use complex functions....Asin(wt)------------>Aexp(iwt) now let's find xp(t)=Aexp(iwt+phi)

OpenStudy (anonymous):

' prime is equivalent to multiply this function by iw

OpenStudy (anonymous):

still with me @SkykhanFalcon

OpenStudy (anonymous):

sorry i had a phone call :/ now i am with u

OpenStudy (anonymous):

what other language can you talk better ? so xp'(t)=iwxp(t) and xp"(t)=-w^2xp(t)

OpenStudy (anonymous):

Turkish but i guess you cannot talk

OpenStudy (anonymous):

No sorry ! then , A(-w^2+iw)exp(iwt+phi)=100exp(i5t)

OpenStudy (anonymous):

it is really hard to understand like that :/

OpenStudy (anonymous):

yeaaah ! I know :( ! It's confusing !

OpenStudy (anonymous):

I'll try to use my equation editor !

OpenStudy (anonymous):

thank u if it works its gonna be perfect for me

OpenStudy (anonymous):

\[A \exp(i \phi)=(1/(- \omega ^2+i \omega))\exp(i(5- \omega )t)\] then \[\omega = 5\] so \[Aexp(i \phi)=1/5(-5+i)\]

OpenStudy (anonymous):

exp(i) is meaning \[e^{i}\] right ?

OpenStudy (anonymous):

so we can choose \[A=1/(5\sqrt{26})\] and phi the argument of that complex number so xp(t)=Asin(omega t+phi) yeaaah right! an BE right back kk !?

OpenStudy (anonymous):

okey

OpenStudy (anonymous):

still heere ?!

OpenStudy (anonymous):

@Neemo yeah i am here

OpenStudy (anonymous):

did you understand or not ?!

OpenStudy (anonymous):

i am lost neemo

OpenStudy (anonymous):

maybe i have to study much more

OpenStudy (anonymous):

Sure ; this is useful for physics right ?!

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

\[\sin(\omega t+\phi)-----> \exp (i \omega t+\phi)\]

OpenStudy (anonymous):

now the operation will be done to exp ... If for physics ; the we got to find a xp \[xp(t)=A \exp (i \omega t+\phi)\]you will calculcate the first deviative · did you get it now ?!

OpenStudy (anonymous):

now the operation will be done to exp ... If for physics ; then we got to find a xp xp(t)=Aexp(iωt+ϕ) you will calculcate the first deviative · did you get it now ?!

OpenStudy (anonymous):

it is very different than i check into the book's answer key

OpenStudy (anonymous):

it is -10/13(5sin5t+cos5t)

OpenStudy (anonymous):

I think it's the same ! we will find xp(t)=Asin(wt+phi) and we will use trig formula's to put sin(wt+phit) that way ! ! But certainly there is something wrong with my answer ! probably a miscalculation !